Grade 9Inorganic Chemistry

50ml of 0.1M NH4OH ,25ml NH4Cl weere used to make a buffer . what is the PH is if Pkbis 4.8?

krishna sanapala
13 Years agoGrade 9

1 Answer

VEERU  YERRAYYA
ApprovedApproved Tutor Answer13 Years ago

The given buffer is a base buffer

[NH4OH] =[(50*0.1)/75]

[NH4Cl] =[(25*2)/75]

From the principle V1M1=V2M2

For a base buffer pOH = pKb + log [Salt/Base]

Therefore pOH 4.8+ log [{(25*2)/75}/{(50*0.1)/75}] = 4.8+ log (10) = 4.8+1 =5.8

Therefore pH = (14-pOH) = (14-5.8) = 8.2